OpenStudy (anonymous):

Write an equation in slope-intercept form of the line that passes through (4, 8) and (0, -2).

OpenStudy (anonymous):

y=5/2x-2 please check cuz that was in my head X(

OpenStudy (anonymous):

First find slope m = (8 - -2)/(4 - 0) = 10/4 = 5/2 Now use point slope y + 2 = (5/2)(x - 0) subtract 2 to both sides y = (5/2)x - 2

OpenStudy (anonymous):

Thnks Guys.

OpenStudy (anonymous):

XD